OpenStudy (anonymous):

Solve x + 3 = 11. A. –8 B. 8 C. 14 D. –14

OpenStudy (anonymous):

Just subtract 3. x + 3 = 11 x = 8

OpenStudy (anonymous):

o ok i see thanks

OpenStudy (anonymous):

Solve 13 – x = 4. A. –9 B. 17 C. 9 D. –17

OpenStudy (anonymous):

13-4=9

OpenStudy (anonymous):

ok that a little easyer thanks alot but i still have a few

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Please try to abide to the code of conduct UditKulka.

OpenStudy (anonymous):

Solve 7x = –21. A. –14 B. –28 C. –3 D. –21

OpenStudy (anonymous):

and 86camaro as well, for that matter

OpenStudy (anonymous):

is it -3

OpenStudy (anonymous):

Yes, how did you arrive to your conclusion?

OpenStudy (anonymous):

i guessed

OpenStudy (anonymous):

it is -3 because you divide -21 by 7